OpenStudy (anonymous):

A pyramid has a volume of 16 and an altitude of 5. What is the area of its base?

OpenStudy (anonymous):

the volume of a pyramid is given by: v=1/3*A*H

OpenStudy (anonymous):

We have : 16=1/3*A*5, how should be or rather: 16=5/3*A, how should we solve this for A

OpenStudy (anonymous):

do you have an idea on how to solve for A??

OpenStudy (anonymous):

not at all

OpenStudy (anonymous):

Well, since we are trying to get A along, but A is mulitplied by 5/3, what process allows us to undo multiplication??

OpenStudy (anonymous):

A = 9.6

OpenStudy (anonymous):

yes